# Half-VAE: An Encoder-Free VAE to Bypass Explicit Inverse Mapping

### v1

@@ 1 Introduction
Due to the integration of variational Bayesian theory, the encoding process in VAE can be regarded as an inference process. When VAE is designed to solve problems such as ICA or disentanglement, this inference process can be equivalent to an inverse mapping. Briefly, when considering the latent variable $\mathbf{Z}$ as independent components, the encoding process in a VAE represents the inverse mapping from the mixed observations to the independent components, i.e., $\mathbf{Z}=f^{-1}(\mathbf{X})$. On the other hand, the decoding process in a VAE represents the remapping of the independent components back into the observed data, i.e., $\mathbf{X}=f(\mathbf{Z})$. In many previous studies, VAE methods for performing disentanglement or ICA typically have a prerequisite that $M\geq N$ (Casale et al., (2018); Wei, (2024)). This is because the encoder struggles to handle situations where the dimensionality $M$ of the input data $\mathbf{X}$ is smaller than the dimensionality $N$ of the latent variable $\mathbf{Z}$ (also known as the underdetermined case (Comon, (1994))). Under underdetermined conditions, the inverse mapping $\mathbf{f}^{-1}$ is theoretically nonexistent (Cardoso, (1998); Bofill and Zibulevsky, (2001); Hyvarinen et al., (2001); Cichocki, (2002)). Therefore, a strategy is needed to bypass the explicit inverse mapping process in the VAE architecture.
@@
This study discards the encoder in the VAE architecture, allowing the posterior distributions of the latent variables to be directly obtained through trainable parameters converging based on the objective function. This encoder-free VAE is referred to as Half-VAE. We use several randomly generated independent and identically distributed (i.i.d.) signals as the independent components and select Gaussian Mixture Models (GMM) (Dempster et al., (1977); McLachlan, (2000); Bishop, (2006); Reynolds et al., (2009)) as the prior distributions, with the GMM parameters set as optimizable, adaptive parameters. In this design, the parameters of the posterior and prior distributions of the latent variables, as well as the decoder’s parameters, are all optimized through the objective function. From the perspective of parameter optimization, a typical VAE requires optimizing the parameters of both the encoder and decoder, whereas Half-VAE directly optimizes the posterior distributions of the latent variables. We find that even without an encoder, the different dimensions of the Half-VAE’s latent variables can still converge to mutually independent i.i.d. sequences. Discarding the encoder is equivalent to discarding the inverse mapping stage, theoretically leaving only the ’forward’ process from the independent components to the observed signals in the Half-VAE. In other words, there is no explicit inverse mapping included in the process of inferring the independent components.
@@
The primary contribution of this study is the introduction of the Half-VAE, an innovative encoder-free VAE architecture. We derive the variational lower bound for the Half-VAE, which serves as its objective function, and validate its effectiveness in solving ICA problems through numerical experiments. Our results demonstrate the feasibility of solving ICA without the need for an explicit inverse mapping process, establishing the practicality of this architecture for such tasks.
@@ 2.1 Derivation
In fact, in the VAE loss function, maximizing the ELBO forces $D_{\text{KL}}(q(\mathbf{Z}|\mathbf{X})\|p(\mathbf{Z}|\mathbf{X}))$ to approach zero, thereby making $q(\mathbf{Z}|\mathbf{X})$ approximate $p(\mathbf{Z}|\mathbf{X})$. Here, $p(\mathbf{Z}|\mathbf{X})$ represents the true posterior of the latent variables, while $q(\mathbf{Z}|\mathbf{X})$ is an additional term introduced for approximation. In typical VAEs, this introduced $q(\mathbf{Z}|\mathbf{X})$ is represented by the encoder, which maps directly from $\mathbf{X}$ to $\mathbf{Z}$. However, in the theoretical derivation of variational inference, the term used to approximate $p(\mathbf{Z}|\mathbf{X})$ does not necessarily have to be $q(\mathbf{Z}|\mathbf{X})$. In this study, $I(\mathbf{Z})$ is directly used to approximate $p(\mathbf{Z}|\mathbf{X})$. Although the explicit encoding process is removed, the core principle of variational inference remains unchanged.
@@ A Gaussian Mixture Model (GMM) is defined as:
where $\lambda$ is used to adjust the weights of each item in the loss functions. The $I(\mathbf{Z})$ directly represented by $\mathbf{Z}_{\mu}$ and $\mathbf{Z}_{\sigma}$. Note that, for the convenience of performing stochastic gradient descent in neural networks, we have reformulated the maximization of the ELBO (see Equation 7) into the minimization of the negative ELBO as loss functions, as shown in Equations 8 and 9.
@@ 4 Discussions and future works
This study proposes a novel VAE architecture, the Half-VAE, and derives its variational lower bound, which also serves as the objective function. Unlike the traditional VAE architecture, the Half-VAE discards the encoder, with the latent variables directly represented by trainable parameters that are optimized through the loss function. Since the Half-VAE does not have the encoding process present in a VAE, it avoids the need for explicit inverse mapping. Our experiments demonstrate that the Half-VAE performs comparably to the traditional VAE in solving ICA problems.
However, several issues require further investigation. At the current stage of research, the Half-VAE still cannot directly achieve ICA under underdetermined conditions. Although the Half-VAE avoids explicit inverse mapping, solving underdetermined problems still requires more assumptions and constraints depending on the various conditions (Comon, (1994); Hyvarinen et al., (2001)). Additionally, we found that the random initialization of parameters in both the Half-VAE and VAE architectures plays a crucial role in how quickly the models converge. Thus, future work will also investigate the design of effective initialization strategies for both Half-VAE and VAE in solving ICA problems.

### v2

@@ 1 Introduction
Due to the integration of variational Bayesian theory, the encoding process in VAE can be regarded as an inference process. When VAE is designed to solve problems such as ICA or disentanglement, this inference process can be equivalent to an inverse mapping. Briefly, when considering the latent variable $\mathbf{Z}$ as independent components, the encoding process in a VAE represents the inverse mapping from the mixed observations to the independent components, i.e., $\mathbf{Z}=f^{-1}(\mathbf{X})$. On the other hand, the decoding process in a VAE represents the remapping of the independent components back into the observed data, i.e., $\mathbf{X}=f(\mathbf{Z})$. Generally, $f$ represents static mixing, which means the mixing process remains constant and does not vary over time or space. In many previous studies, methods using Variational Autoencoders (VAEs) for performing disentanglement or Independent Component Analysis (ICA) typically have a prerequisite that the dimensionality $M$ of the input data $\mathbf{X}$ is at least as large as the dimensionality $N$ of the latent variable $\mathbf{Z}$ Casale et al., (2018); Wei, (2024). This requirement, where $M\geq N$, is crucial because the encoder struggles to handle underdetermined situations, where $M$ is smaller than $N$. In such underdetermined cases, the theoretical inverse mapping ${f}^{-1}$ does not exist Comon, (1994); Cardoso, (1998); Bofill and Zibulevsky, (2001); Hyvarinen et al., (2001); Cichocki, (2002).Therefore, a strategy is needed to bypass the explicit inverse mapping process in the VAE architecture.
There has been some research exploring tasks by removing the encoder from the encoder-decoder architecture. Bond-Taylor and Willcocks, (2020) investigated the generalization capability and reconstruction quality of VAEs after removing the encoder. Shi et al., (2020) designed an encoder-free network to address 3D shape description and retrieval, primarily using Maximum Likelihood Estimation for optimizing neural networks instead of variational inference. Cervantes et al., (2022) proposed Implicit Neural Representations to generate variable-length human motion sequences, employing an encoder-free architecture as well. These studies demonstrate the viability of encoder-free architectures in certain domains. However, no research has yet explored using an encoder-free architecture to bypass the explicit inverse process in solving ICA problems. Considering that the inverse mapping ${f}^{-1}$, represented by the encoding process, does not exist under underdetermined conditions, one possible approach is to circumvent this challenging mapping by omitting the encoder.
@@
This study discards the encoder in the VAE architecture, allowing the posterior distributions of the latent variables to be directly obtained through trainable parameters converging based on the objective function. This encoder-free VAE is referred to as Half-VAE. We use several randomly generated independent and identically distributed (i.i.d.) signals as the independent components and select Gaussian Mixture Models (GMM) (Dempster et al., (1977); McLachlan, (2000); Bishop, (2006); Reynolds et al., (2009)) as the prior distributions, with the GMM parameters set as optimizable, adaptive parameters. In this design, the parameters of the posterior and prior distributions of the latent variables, as well as the decoder’s parameters, are all optimized through the objective function. From the perspective of parameter optimization, a typical VAE requires optimizing the parameters of both the encoder and decoder, whereas Half-VAE directly optimizes the posterior distributions of the latent variables. We find that even without an encoder, the different dimensions of the Half-VAE’s latent variables can still converge to mutually independent i.i.d. sequences. Discarding the encoder is equivalent to discarding the explicit inverse mapping stage, leaving only the ’forward’ process from the independent components to the observed signals.
@@
The primary contribution of this study is the introduction of the Half-VAE, an encoder-free VAE architecture designed to avoid explicit inverse mapping in solving the ICA problem. This research derives the variational lower bound of the Half-VAE and uses it as the objective function. Through simulation experiments, the effectiveness of the Half-VAE in addressing the ICA problem has been validated, laying a foundation for future research.
@@ 2.1 Derivation
In fact, in the VAE loss function, maximizing the ELBO forces $D_{\text{KL}}(q(\mathbf{Z}|\mathbf{X})\|p(\mathbf{Z}|\mathbf{X}))$ to approach zero, thereby making $q(\mathbf{Z}|\mathbf{X})$ approximate $p(\mathbf{Z}|\mathbf{X})$. Here, $p(\mathbf{Z}|\mathbf{X})$ represents the true posterior of the latent variables, while $q(\mathbf{Z}|\mathbf{X})$ is an additional term introduced for approximation. In typical VAEs, this introduced $q(\mathbf{Z}|\mathbf{X})$ is represented by the encoding process, which maps directly from $\mathbf{X}$ to $\mathbf{Z}$. However, in the theoretical derivation of variational inference, the term used to approximate $p(\mathbf{Z}|\mathbf{X})$ does not necessarily have to be $q(\mathbf{Z}|\mathbf{X})$. In this study, $I(\mathbf{Z})$ is directly used to approximate $p(\mathbf{Z}|\mathbf{X})$. Although the explicit encoding process is removed, the core principle of variational inference remains unchanged.
@@ A Gaussian Mixture Model (GMM) is defined as:
where $\lambda$ is used to adjust the weights of each item in the loss functions. The $I(\mathbf{Z})$ directly represented by $\mathbf{Z}_{\mu}$ and $\mathbf{Z}_{\sigma}$. Note that, for the convenience of performing stochastic gradient descent in neural networks, we have reformulated the maximization of the ELBO (see Equation 7) into the minimization of the negative ELBO as loss functions, as shown in Equations 9 and 10. It is important to emphasize that each independent component $\mathbf{Z}_{i}$ has its corresponding prior $\text{GMM}_{\bm{\Psi}_{i}}$. Since the components $\mathbf{Z}_{i}$ are mutually independent, the total KL divergence term in Equations (9) and (10) can be further expressed as the sum of the KL divergences for each independent component:
$\begin{split}D_{\text{KL}}\left[q_{\bm{\phi}}(\mathbf{Z}|\mathbf{X})\|\text{GMM}_{\bm{\Psi}}(\mathbf{Z})\right]=\\ \sum_{i=1}^{N}D_{\text{KL}}\left(q_{\bm{\phi}}(\mathbf{Z}_{i}|\mathbf{X})\|\text{GMM}_{\bm{\Psi}_{i}}(\mathbf{Z}_{i})\right),\end{split}$ (11)
$\begin{split}D_{\text{KL}}\left(I(\mathbf{Z})\|\text{GMM}_{\bm{\Psi}}(\mathbf{Z})\right)=\sum_{i=1}^{N}D_{\text{KL}}\left(I(\mathbf{Z}_{i})\|\text{GMM}_{\bm{\Psi}_{i}}(\mathbf{Z}_{i})\right).\end{split}$ (12)
It is worth noting that the GMM is just one of many available prior distributions. For the Half-VAE, different types of priors can be considered when dealing with signals of varying characteristics.
@@ 4 Discussions and future works
This study introduces the Half-VAE, designed to solve ICA problems without explicit inverse mapping. Unlike the traditional VAE architecture, the Half-VAE discards the encoder, with the latent variables directly represented by trainable parameters optimized through the loss function. The research shows that with appropriate prior settings for the latent variables, the Half-VAE is capable of solving ICA problems. In fact, the performance of the Half-VAE in our numerical examples slightly surpasses that of a VAE with an encoder under the same prior conditions.
However, several issues require further investigation. At the current stage of research, the Half-VAE still cannot directly solve ICA problems under underdetermined conditions. Although the Half-VAE avoids explicit inverse mapping, solving underdetermined problems still requires more assumptions and constraints depending on the various conditions (Comon, (1994); Hyvarinen et al., (2001)). Additionally, we found that the random initialization of parameters in both the Half-VAE and VAE architectures plays a crucial role in how quickly the models converge. Thus, future work will also investigate the design of effective initialization strategies for both Half-VAE and VAE in solving ICA problems. Finally, beyond the GMM, there are many other suitable priors that can be employed. The selection and comparison of appropriate priors for signals with different characteristics will be addressed and explored in future research.